ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 15 Sheth, Dilip N., Sir Parashurambhau College, Pune, India Completion of a unimodular vector to an invertible matrix A. Suslin, in his doctoral thesis, had shown that if the unimodular vector (x_{0}, x_{1}, x_{2}, ⋯, x_{n}) ∈ Um_{n+1}(R) (R a commutative ring with unity), then (x_{0}, x_{1}, x_{2}^{2}, ⋯, x_{n}^{n}) can always be completed to an invertible matrix. In this paper we give an algorithm to construct such an invertible matrix concretely. Reference:

1. A. A. Suslin, On Stably free modules, Math. USSR Sb. 31(1977),479-